it's says: solve by substitution. x-3y=-8 and y=x+2. I alreay know the answer, its (1, 3) but everytime i work it i come up with something completely different.
Answer by rwm(914)

Where is your work?
x-3y=-8
y=x+2
x-3(x+2)=-8
x-3x-6=-8
-2x=-2
x=1
y=x+2